FELION: A CRYOGENIC ION TRAP APPARATUS FOR SPECTROSCOPIC STUDIES WITH FELIX

\underline{S. BRÜNKEN}, L. KLUGE, S. FANGHÄNEL, A. POTAPOV, O. ASVANY, and S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; J. OOMENS, B. REDLICH, Radboud University Nijmegen, Institute for Molecules and Materials, FELIX Facility, 6525 ED Nijmegen, Netherlands; A. STOFFELS, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany, and Radboud University Nijmegen, Institute for Molecules and Materials, FELIX Facility, 6525 ED Nijmegen, Netherlands.

The combination of ion trapping techniques with sensitive action spectroscopy schemes has been developed in recent years as a powerful tool to obtain spectra of gas-phase molecular ions from the UV to the (F)IR spectral regions.^{a} Here we report on the status of a specifically designed, dedicated cryogenic 22-pole ion trap apparatus (FELion), developed and built in Cologne, Germany, and intended to be installed permanently at the ”Free-Electron Lasers for Infrared eXperiments” (FELIX) facility in Nijmegen, the Netherlands. This instrument will allow to record gas-phase IR and FIR spectra of mass-selected, internally cold molecular ions at temperatures in the range 4 − 300 K. By the use of diverse ionization methods, e.g. electron impact and electrospray ionization, a multitude of molecular ions can be generated and stored in the trap, including astrophysically relevant species ranging in size from the three-atomic H_{3}^{+} up to large polycyclic aromatic hydrocarbon (PAH) ions, but also biomolecular ions like amino acids, peptides, or nucleobases. In combination with the powerful (F)IR radiation of the free electron lasers FELIX-1 and -2 (60−2500 cm^{−1}) and FLARE^{b} (6−100 cm^{−1}) at the FELIX facility, a variety of action spectroscopy schemes can be employed to study the ro-vibrational spectra of the stored ions, such as IR multiphoton dissociation, (F)IR/UV double resonance spectroscopy, or the method of laser induced reactions (LIR). In this talk we will give a detailed account of the experimental setup and present the first results obtained with the new apparatus.

ae.g., S. Schlemmer, E. Lescop, J. von Richthofen, D. Gerlich, and M. A. Smith, J. Chem. Phys. 117, 2068 (2002); J. Oomens, B. G. Sartakov, G. Meijer, and G. van Helden, Int. J. Mass Spectrom. 254, 1 (2006); T. R. Rizzo, J. A. Stearns, and O. V. Boyarkin, International Reviews in Physical Chemistry 28, 481 (2009)

bFLARE: Free-electron Laser for Advanced spectroscopy and high-Resolution Experiments